OpenStudy (anonymous):

Find the slope and the y-intercept of the line. y= -3/2x+4 slope: y-intercept:

OpenStudy (serenity74):

y = mx + b b = y-intercept m = slope So what would be the Slope and y intercept in this case?

OpenStudy (mathmale):

Strongly suggest that you present those fractions inside parentheses. y= -3/2x+4 should be (-3/2)x + 4. (-3/2) is the coefficient (multiplier) of the x term, and the 4 is the constant term. Does this help you to identify the slope? The y-intercept?

OpenStudy (mathmale):

Compare (-3/2)x + 4 to y = m*x + b and read off the slope and y-intercept.